if a load of 200° is moved up an inclined plane at an angle 30° to the horizontal using a machine of efficiency of 80%.calculate the effort applied


1
Expert's answer
2021-06-02T09:35:33-0400
"0.8=\frac{200l\sin{30}}{Fl}\\0.8=\frac{100}{F}\\F=125\ N"